# Charged pions asymmetry from the decay of light neutral axial mesons due to interference

### v1

@@ 1 Introduction
Three meson interactions based in QCD- one loop quark polarization have been calculated by considering, for example, the Global Color Model (GCM) or the Nambu-Jona-Lasinio (NJL)-model [20, 21, 22, 6, 7]. Dynamical Chiral Symmetry Breaking (DChSB) endows constituent quarks with large quark masses and this makes possible large quark mass expansions of the quark determinant. Flavor-Dirac quark currents provide suitable states to introduce meson multiplets, already coping with the constituent quark model, what can be implemented by means of the auxiliary field method. However, there are longstanding controversial results concerning light scalar and axial mesons structures, for which non-quark-antiquark states may contribute considerably [23, 25, 24]. Nevertheless, along this work, it will be assumed that an axial meson nonet and a light scalar meson nonet may be defined, although the physical states may be a result of superposition with different non-quark-antiquark components. In particular, the following scalar and axial mesons will be considered as part of flavor U(3) nonets: the isotriplets scalar $a_{0}(980)$ and axial $a_{1}(1260)$, and isoscalar axial mesons $f_{1}(1285)$ and $f_{1S}(1420)$ [25]. The production of axial $f_{1}(1285)$ and $f_{1S}(1420)$ has been investigated in proton-proton collisions [26, 27] and it may involve different types of processes. In the present work, the decay of these two axial mesons into charged pions and another meson - either a charged $\rho(770)$ or charged scalar $a_{0}(980)$, will be investigated with three-meson vertices.
@@
Therefore, in this work, some specific three light meson leading interactions are re-derived, along the lines of [7] and used to investigate neutral axial meson decays into charged pions. These flavor U(3) interactions describe several channels, among which only processes involving mesons composed exclusively by up and down quarks in the final states will be considered. The channels with strange mesons will necessarily involve states for which there are still further problems to define quark-antiquark mesons, possibly in possible scalar and axial meson nonets, and they will be investigated in a separate work. Furthermore, the point-like limit of the three-meson interactions, and the corresponding zero momentum limit of the form factors, will be considered. This will provide an upper bound for the resulting effects. The following decays of neutral axial mesons ($A=f_{1}(1285)$ and $f_{1S}(1420)$): $A\to\pi^{\mp}+a_{0}^{\pm}$ (processes ASP) and $A\to\pi^{\mp}+\rho^{\pm}$ (processes VPA). To complement the description, the decay $\rho^{\pm}\to\pi^{\pm}+\pi^{0}$ with large relative branching ratio - will also be presented. The work is organized as follows. In the next section, the flavor U(3) quark determinant with background (constituent) quark currents is considered in the limit of local quark currents that give rise to the corresponding light meson multiplets by means of the auxiliary field method (AFM). All the local couplings are resolved by means of a zeroth order derivative expansion in the limit of zero momentum exchange, leading to coupling constants along the lines presented in [7]. In section (2.1) three-meson interactions for two channels of neutral axial meson decays, $f_{1}(1285)$ and $f_{1S}(1420)$, into charged pions and rhos $\rho^{\pm}(770)$ or into charged pions and scalars $a_{0}^{\pm}(980)$ will be presented, named respectively as VPA and ASP. For the channel VPA a next leading interaction will be also considered as shown in section (2.1.1). In section (2.2) the leading interactions describing the decay $\rho^{\pm}\to\pi^{\pm}\pi^{0}$, VPP, are presented. In sections (2.3) and (2.4) the renormalization conditions fixing free parameters of the AFM and the definition of meson states, considered along the work, are presented. In section (3) the different ratios $\Gamma_{\pi^{+}}/\Gamma_{\pi^{-}}$ for each of the decays are calculated by including meson mixings. Numerical results are exhibited in (4). In the last section there is a summary.
@@ 2 Quark determinant with background quark currents: and meson local interactions
where $\lambda_{i}$ are the flavor Gell Mann matrices, flavor indices of the adjoint representation are $i,j=0...(N_{f}-1)^{2}$, $K_{0}=2\alpha_{g}^{2}/9$ being $2/9$ from a Fierz transformation and $\alpha_{g}$ the running quark-gluon coupling constant. The quark currents were defined respectively as $J_{S}^{i}=J_{S}^{i}(x,y)=(\bar{\psi}\lambda^{i}\psi)$, $J_{PS}^{i}=J_{PS}^{i}(x,y)=(\bar{\psi}i\gamma_{5}\lambda^{i}\psi)$, $J_{V,i}^{\mu}=J_{V,i}^{\mu}(x,y)=(\bar{\psi}\gamma^{\mu}\lambda^{i}\psi)$ and $J_{A,i}^{\mu}=J_{A,i}^{\mu}(x,y)=(\bar{\psi}\gamma^{\mu}\gamma_{5}\lambda^{i}\psi)$. Note that, in Eq. (1), the background quark currents are dressed by components of the effective gluon propagator, and they can be considered as constituent quark currents. The large quark mass expansion done below is also compatible with a large gluon effective mass expansion since all the coefficients $a_{\phi}$ in Eq. (1) contain components of the gluon propagator making the overall quantities $a_{\phi}J_{\phi}$ with relative small strength. Although the resulting model is non-local, the large quark mass (and to some extent the large effective gluon mass) regime can be considered suitable to take the local limit of such quark currents. In this work, the indices of the currents will be treated such as it does not matter whether they stand as superscripts or subscripts. The functions $R(x-y)$ and $R^{\mu\nu}(x-y)$ are expressed in terms of the longitudinal and transverse components of the gluon propagator and they can be written in momentum space as:
@@
There are several ways to treat this determinant and a direct large quark (and gluon) effective mass(es) expansion, will be performed.
The auxiliary field method is extremely useful to implement quark-antiquark states from the quark currents. It produces quark model meson multiplets naturally based in flavor symmetry. However, instead of introducing the non-local auxiliary (meson) fields, the local limit of the above quark currents will be considered and local meson fields for each of the channels (scalar, pseudoscalar, vector and axial) will be introduced by means of a functional delta functions [32] that can be written as
@@ 2.1 VPA and ASP vertices: neutral axial mesons decays
where $T^{ijk}=2(d_{ijk}+if_{ijk})$. By defining the function $R^{\eta}(k)$ to be either the longitudinal $\eta=L$ or the transverse $\eta=T$ part of the gluon propagator, the following flavor dependent coupling constants were defined, for the limit of zero momentum exchange:
@@
After having defined the coupling constants at the zero momentum exchange, i.e. neglecting the momentum dependence, all the energy dependencies of the couplings in Eq. (6) will also be neglected by fixing the longitudinal vector and axial meson polarizations, such as $\mu=0$, and fixing the meson energy as its rest energy, i.e. $E_{\phi}=M_{\phi}$.
@@ 2.2 VPP: Charged rho meson decay into two pions
There are several ways to treat the quark determinant. Strictly, the expansion shown above does not describe several channels, such as the charged vector meson decays into two pions, $\rho^{\pm}\to\pi^{\pm}\pi^{0}$. To do that, let us define the following quantity in the determinant
@@ 2.4 Meson states
The description of light scalars structure has controversial results. However, it will be considered that the isotriplet states $S_{1,2,3}$ are defined as quark-antiquark mesons $a_{0}(980)$ [25]. The axial nonet is not well established neither [25, 24] and to avoid further uncertainties in the calculation, the analysis will be restricted to final states containing only mesons with up and down quarks. The following states will be considered: [25, 5]:
@@ 3 Asymmetries in the decay of light axial mesons into charged pions
Some of these neutral axial mesons decays, as the initial and final states, have already been measured experimentally and others haven’t [25]:
(observed) $\;\;\;\;A_{1}(1260)\to\rho(770)\pi,\;\;\;\;f_{1}(1285)\sim A_{0}\to\rho\pi,\;\;\;\;\;f_{1}(1285)\to a_{0}(980)\pi,$ (23)
(not observed) $\;\;\;\;A_{1}(1260)\to a_{0}^{0}(980)\pi.\;\;\;\;f_{1S}(1420)\sim A_{8}\to\rho\pi,\;\;\;\;f_{1S}(1420)\to a_{0}(980)\pi.$
Some of these decay may respect the OZI sum rule and others do not [24]. The decays considered for the $f_{1S}(1420)$ meson have not been observed experimentally and this might be due to several reasons such as very small branching ratio, difficulties with final state interactions or even that its structure does not correspond exactly to the state $A_{8}$ of a quark-antiquark nonet. However, it might also happen that state $A_{8}$ of the possible axial meson nonet can undergo strong mixings with the states $A_{0}$ and $A_{3}$ as discussed below and these mixings might prevent a clear identification of its structure at the example of the $\eta-\eta^{\prime}$ mixing of the pseudoscalar mesons multiplet. Therefore, the channels presented above will be investigated below, at least, as a possible scenario. Besides that, there are decays of the charged rho $\rho$:
@@ 3.1 VPA-A decay
${\cal L}_{A\rho^{\mp}\pi^{\pm}}^{f}=12\;i\;\left[G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right]\left[\pi_{3}(\rho^{-}A_{1}^{+}-\rho^{+}A_{1}^{-})+A_{1}^{3}(\rho^{+}\pi^{-}-\rho^{-}\pi^{+})+\rho_{3}(A_{1}^{-}\pi^{+}-A_{1}^{+}\pi^{-})\right],$ (26)
@@ 3.1.1 Meson mixing
${\cal V}_{A\rho^{+}\pi^{-}}^{f}$ ::: $12\;i\;\left[G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right]\;D_{A_{3}}(P^{2})\;\left(A_{1}^{8}\;\bar{G}_{38}+A_{1}^{0}\;\bar{G}_{03}\right),$
${\cal V}_{A\rho^{-}\pi^{+}}^{f}$ ::: $-12\;i\;\left[G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right]\;D_{A_{3}}(P^{2})\;\left(A_{1}^{8}\;\bar{G}_{38}+A_{1}^{0}\;\bar{G}_{03}\right).$ (34)
@@ 3.3.1 VPA followed by VPP
where $J=0,8,$, $D_{\rho}(P_{\rho}^{2})$ and $D_{\eta}(P_{\eta}^{2})$ are respectively the rho and eta propagators with corresponding momenta according to Figs. (3) and (2).

### v3

# Charged pions asymmetry due to interference from the decay of light neutral axial mesons

@@ 1 Introduction
Three meson interactions based in QCD- one loop quark polarization have been calculated by considering, for example, the Global Color Model (GCM) or the Nambu-Jona-Lasinio (NJL)-model [20, 21, 22, 6, 7]. Dynamical Chiral Symmetry Breaking (DChSB) endows constituent quarks with large quark masses and this makes possible large quark mass expansions of the quark determinant. Flavor-Dirac quark currents provide suitable states to introduce meson multiplets, already coping with the constituent quark model, what can be implemented by means of the auxiliary field method. However, there are longstanding controversial results concerning light scalar and axial mesons structures, for which non-quark-antiquark states may contribute considerably [23, 25, 24]. Nevertheless, along this work, it will be assumed that an axial meson nonet and a light scalar meson nonet may be defined, although the physical states may be a result of superposition with different non-quark-antiquark components. In particular, the following scalar and axial mesons will be considered as part of flavor U(3) nonets: the isotriplets scalar $a_{0}(980)$ and axial $a_{1}(1260)$, and isoscalar axial mesons $f_{1}(1285)$ and $f_{1S}(1420)$ [25]. The production of axial $f_{1}(1285)$ and $f_{1S}(1420)$ has been investigated in proton-proton collisions [26, 27, 28] and it may involve different types of processes. In the present work, the decay of these two axial mesons into charged pions and another meson - either a charged $\rho(770)$ or charged scalar $a_{0}(980)$, will be investigated with three-meson vertices.
@@
Therefore, in this work, some specific three light meson leading interactions are re-derived, along the lines of [7] and used to investigate neutral axial meson decays into charged pions. These flavor U(3) interactions describe several channels, among which only processes involving mesons composed exclusively by up and down quarks in the final states will be considered. The channels with strange mesons will necessarily involve states for which there are still further problems to define quark-antiquark mesons, possibly in possible scalar and axial meson nonets, and they will be investigated in a separate work. Furthermore, the point-like limit of the three-meson interactions, and the corresponding zero momentum limit of the form factors, will be considered. This will provide an upper bound for the resulting effects. The following decays of neutral axial mesons will be considered ($A=f_{1}(1285)$ and $f_{1S}(1420)$): $A\to\pi^{\mp}+a_{0}^{\pm}$ (processes containing one axial, one scalar and one pseudoscalar mesons will be labeled by ASP) and $A\to\pi^{\mp}+\rho^{\pm}$ (processes containing one axial, one vector and one pseudoscalar mesons will be labeled by VPA). To complement the description, the decay $\rho^{\pm}\to\pi^{\pm}+\pi^{0}$ with large relative branching ratio - will also be presented. The work is organized as follows. In the next section, the flavor U(3) quark determinant with background (constituent) quark currents is considered in the limit of local quark currents that give rise to the corresponding light meson multiplets by means of the auxiliary field method (AFM). All the local couplings are resolved by means of a zeroth order derivative expansion in the limit of zero momentum exchange, leading to coupling constants along the lines presented in [7]. In section (2.1) three-meson interactions for two channels of neutral axial meson decays, $f_{1}(1285)$ and $f_{1S}(1420)$, into charged pions and rhos $\rho^{\pm}(770)$ or into charged pions and scalars $a_{0}^{\pm}(980)$ will be presented, named respectively as VPA and ASP. For the channel VPA a next leading interaction will be also considered as shown in section (2.1.1). In section (2.2) the leading interactions describing the decay $\rho^{\pm}\to\pi^{\pm}\pi^{0}$, VPP, are presented. In sections (2.3) and (2.4) the renormalization conditions fixing free parameters of the AFM and the definition of meson states, considered along the work, are presented. In section (3) the different ratios $\Gamma_{\pi^{+}}/\Gamma_{\pi^{-}}$ for each of the decays are calculated by including meson mixings. Numerical results are exhibited in (4). In the last section there is a summary.
@@ 2 Quark determinant with background quark currents: and meson local interactions
where $\lambda_{i}$ are the flavor Gell Mann matrices, flavor indices of the adjoint representation are $i,j=0...(N_{f}-1)^{2}$, $K_{0}=2\alpha_{g}^{2}/9$ being $2/9$ from a Fierz transformation and $\alpha_{g}$ the running quark-gluon coupling constant. The quark currents were defined respectively as $J_{S}^{i}=J_{S}^{i}(x,y)=(\bar{\psi}\lambda^{i}\psi)$, $J_{PS}^{i}=J_{PS}^{i}(x,y)=(\bar{\psi}i\gamma_{5}\lambda^{i}\psi)$, $J_{V,i}^{\mu}=J_{V,i}^{\mu}(x,y)=(\bar{\psi}\gamma^{\mu}\lambda^{i}\psi)$ and $J_{A,i}^{\mu}=J_{A,i}^{\mu}(x,y)=(\bar{\psi}\gamma^{\mu}\gamma_{5}\lambda^{i}\psi)$. Note that, in Eq. (1), the background quark currents are dressed by components of the effective gluon propagator $R(x-y)$ and $R^{\mu\nu}(x-y)$ that are defined below, and they can be considered as constituent quark currents. The large quark mass expansion done below is also compatible with a large gluon effective mass expansion since all the coefficients $a_{\phi}$ in Eq. (1) contain components of the gluon propagator making the overall quantities $a_{\phi}J_{\phi}$ with relative small strength. Although the resulting model is non-local, the large quark mass (and to some extent the large effective gluon mass) regime can be considered suitable to take the local limit of such quark currents. In this work, the indices of the currents will be treated such as it does not matter whether they stand as superscripts or subscripts. The functions $R(x-y)$ and $R^{\mu\nu}(x-y)$ are expressed in terms of the longitudinal and transverse components of the gluon propagator and they can be written in momentum space as:
@@
[…]
The auxiliary field method is extremely useful to implement quark-antiquark states from the quark currents. It produces quark model meson multiplets naturally based in flavor symmetry. However, instead of introducing the non-local auxiliary (meson) fields, the local limit of the above quark currents will be considered and local meson fields for each of the channels (scalar, pseudoscalar, vector and axial) will be introduced by means of functional delta functions [31] that can be written as
@@ 2.1 VPA and ASP vertices: neutral axial mesons decays
where $T^{ijk}=2(d_{ijk}+if_{ijk})$, in terms of the flavor group symmetric and anti-symmetric structure constants respectively. By defining the function $R^{\eta}(k)$ to be either the longitudinal $\eta=L$ or the transverse $\eta=T$ part of the gluon propagator, the following flavor dependent coupling constants were defined, for the limit of zero momentum exchange:
@@
It is interesting to note that, in general, the coupling constants for a particular two meson decay channel are the same for both types of interactions with symmetric and antisymmetric flavor structure constants $d_{ijk}$ and $f_{ijk}$ This means that the specific momentum dependence of such form factors, defined by Eqs. (2.1) and in other equations below, should nearly factorize in the ratios of decay rates explored below. Therefore, these momentum dependencies should not introduce considerable changes in the resulting ratios of meson decay, and the coupling constants will be considered for numerical calculations instead of the form factors. Furthermore, this issue must be related to the problem of fixing the vector and axial meson polarizations for all the corresponding momentum dependent vertices. The inclusion of the full momentum dependent form factors and their behavior, conjugated to the choice of the vector/axial meson polarization, as discussed after Eq. (3.1) and below, will be developed in another work. After having defined the coupling constants at the zero momentum exchange, i.e. neglecting the momentum dependence, all the energy dependencies of the couplings in Eq. (6) will also be neglected by fixing the longitudinal vector and axial meson polarizations, such as $\mu=0$, and fixing the meson energy as its rest energy, i.e. $E_{\phi}=M_{\phi}$.
@@ 2.2 VPP: Charged rho meson decay into two pions
In this section we develop vertices with one vector meson and two pseudoscalar mesons, labeled by VPP. There are several ways to treat the quark determinant. Strictly, the expansion shown above does not describe several channels, such as the charged vector meson decays into two pions, $\rho^{\pm}\to\pi^{\pm}\pi^{0}$. To do that, let us define the following quantity in the determinant
@@ 2.4 Meson states
[…]
@@ 3 Asymmetries in the decay of light axial mesons into charged pions
Most of the axial meson decays considered in this work have already been measured experimentally. The only decays for which no record was found in the PDG Table [25] is $a_{1}(1260)\to a_{0}^{0}(980)\pi$ and $f_{1S}(1420)\to\rho\pi$. The other processes, showed in figures (1), (2) and (3) are the following [25]:
[…]
Some of these decay may respect the OZI sum rule or G-parity conservation and others do not [24]. Besides that, neutral axial meson mixings might prevent a clear identification of their structure at the example of the $\eta-\eta^{\prime}$ mixing of the pseudoscalar mesons multiplet [43, 44]. Therefore, the channels presented above will be investigated below, at least, as a possible scenario. Besides that, there are decays of the charged rho $\rho$:
@@ 3.1 VPA-A decay
${\cal L}_{A\rho^{\mp}\pi^{\pm}}^{f}=12\;i\;\left[G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right]\left[\pi_{3}(\rho^{-}a_{1}^{+}-\rho^{+}a_{1}^{-})+a_{1}^{0}(\rho^{+}\pi^{-}-\rho^{-}\pi^{+})+\rho_{3}(a_{1}^{-}\pi^{+}-a_{1}^{+}\pi^{-})\right],$ (26)
@@ 3.1.1 Meson mixing
${\cal V}_{A\rho^{+}\pi^{-}}^{f}$ ::: $12\;i\;\left[G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right]\;D_{A_{3}}(P^{2})\;\left(A_{1}^{8}\;\bar{G}_{38}+A^{0}_{1}\;\bar{G}_{03}\right),$
${\cal V}_{A\rho^{-}\pi^{+}}^{f}$ ::: $-12\;i\;\left[G_{sb1}^{R}+E_{\pi}E_{A}\;G_{sva,sb}^{R}\right]\;D_{A_{3}}(P^{2})\;\left(A_{1}^{8}\;\bar{G}_{38}+A^{0}_{1}\;\bar{G}_{03}\right).$ (34)
@@ 3.3.1 VPA followed by VPP
where $J=0,8$, $D_{\rho}(P_{\rho}^{2})$ and $D_{\eta}(P_{\eta}^{2})$ are respectively the rho and eta propagators with corresponding momenta according to Figs. (3) and (2).